OpenStudy (jade27p):

What is the slope of the line between (3, −4) and (−2, 1)?

OpenStudy (jade27p):

@.Sam. can you help me please

OpenStudy (jadeishere):

Okay, do you know the slope formula? \[m = \frac{ y _{2} - y _{1} }{ x _{2}- x _{1} }\]

OpenStudy (keythemusician):



OpenStudy (jade27p):

yes

OpenStudy (jadeishere):

Basically, it's a 'plug-n-chug' formula used to find the slope. :) So plug in the y and x values and solve from there :P

OpenStudy (jade27p):

.6

OpenStudy (jadeishere):

So when you plug in the values, it'll look like this; \[m = \frac{ 1-(-4) }{ -2-3}\] 1 - (-4) Since two negatives will equal a positive, it's 5 -2 - 3 Add/subtract these two numbers, and you'll get -5 \[\frac{ 5 }{ -5 } = m\]

OpenStudy (jade27p):

do i divide that

OpenStudy (jadeishere):

Yes :)

OpenStudy (jade27p):

-1

OpenStudy (jadeishere):

Yes, very good! :D

OpenStudy (jade27p):

thank you so much, can you help with one more

OpenStudy (jadeishere):

Of course :)

OpenStudy (jade27p):

im going to make another question to give you another medal

OpenStudy (jadeishere):

Oh,. that's not necessary :P I'm just here to help